OpenStudy (anonymous):

How do I determine if these two lines are perpendicular? y=2x-9 3y=5-x

OpenStudy (anonymous):

First set em up as y=mx+b

OpenStudy (anonymous):

change the bottom one to y = mx+b form so 3y = 5-x y = -1/3x + 5/3 compare the m's in each equation. the m in the first one is 2, and the m in the second one is -1/3 for the lines to be perpendicular, the slopes must be negative inverses of each other...the negative inverse of 2 is -1/2, which is NOT -1/3 so they aren't.

OpenStudy (anonymous):

Thank you for the explanation.

OpenStudy (anonymous):

no problem!